Wow! If this actually works, Im very impressed by this free algebra/tutoring resource. I have several different problems, But would appreciate it if I could have just one of each type of them solved and explained to me how it was done in depth.
Solve Each System Of Equations Using Graphing
1.--y=-6
----4x+y=2




Solve Each System Of Equations By Substitutio
1.x=2y
4x+2y=15



2.2x+7y=3
x=1-4y





3.3y=12
x=y+8





Solve Each System Using Elimination
5x+2y=7
-2x+2y=-14

Answer by venugopalramana(3286)

it is assumed as follows
-y=-6...so y=(-6)/(-1)=6/1=6...put this value in next eqn.
-4x+y=2..we get ..-4x+6=2
-4x=2-6=-4
x=(-4)/(-4)=1
hence x=1 and y=6
******************
x=2y...put this next eqn.
4x+2y=15...we get 4(2y)+2y=15
8y+2y=15
10y=15
y=15/10=3/2
x=2y=2*3/2=3
hence x=3 and y=3/2
*************
2x+7y=3
x=1-4y...you can do it as in the above problem by putting x=1-4y in 2x+7y=3...
***************************************************************************
3y=12
x=y+8...this is same as the first problem which is shown above.you follow the same method...
*********************************************
Solve Each System Using Elimination
5x+2y=7............(1)
-2x+2y=-14........(2)
subtract eqn.(2) from eqn.(1)
5x+2y-(-2x+2y)=7-(-14)
5x+2y+2x-2y=7+14=21
7x=21
x=21/7=3
putting in eqn.(1),we get
5*3+2y=7
2y=7-15=-8
y=-8/2=-4
hence x=3 and y=-4